OpenStudy (anonymous):

simplify the expression: (5x+1/5)^2

OpenStudy (anonymous):



OpenStudy (anonymous):

Have you done the "FOIL" (first, outer, inner, last) method?

OpenStudy (anonymous):

i just learned today but i'm confused

OpenStudy (anonymous):

Ok, you write the expression like: (5x + 1/5) (5x + 1/5) Then multiply the two FIRST terms together... (5x)(5x) = 25x^2 multiply the two OUTER terms together (the first and last terms): (5x)(1/5) = x then multiply the middle or INNER terms: (1/5)(5x) = x Then multiply the two LAST terms (1/5)(1/5) = (1/25) and add it all up

OpenStudy (anonymous):

(5x + 1/5)(5x + 1/5) 25x^2 + x + x + (1/25) then simplify: 25x^2 + 2x + (1/25) \[25^{2}+2x+\frac{ 1 }{ 25 }\]

OpenStudy (anonymous):

i understand now ^.^ thanks!

OpenStudy (anonymous):

Squaring something that has two parts like (a+b) is just multiplying (a+b)(a+b) and the FOIL approach just helps you make sure you distribute everything properly across the parenthesis when you multiply. you already know that a(a+b) would be a^2 + ab because of distribution and b(a+b) would be ab + b^2 the same way

OpenStudy (anonymous):

Great :) Glad to help!